OpenStudy (anonymous):

6z+2=z-4 Solve for z and check solution please

OpenStudy (anonymous):

6z + 2 = z - 4 6z + 2 - 2 = z -4 -2 6z = z - 6 6z -z = z - z - 6 5z = -6 z = -6 / 5 Insert it into the top equation.

sam (.sam.):

\[\begin{array}{l} \text{Add }(-z-2)\text{ \to both sides:} \\ 5 z=-6 \\ \text{Divide both sides by }5: \\ z=-\frac{6}{5} \\\end{array}\]

OpenStudy (anonymous):

thanks Sam